OpenStudy (anonymous):

integral 1/(sqroot(e^(2x)-1)) dx solve using trig substitution. I can solve it using other methods, but I can't using trig substitution...

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{1}{\sqrt{e^{2x}-1}}dx\quad=\quad \int\limits \frac{1}{\sqrt{\left(e^x\right)^2-1}}dx\]Let:\[\Large\bf\sf e^x=\sec \theta\]

OpenStudy (anonymous):

multiply both top and bottom by e^x, and then let u = e^x.

zepdrix (zepdrix):

Do you understand how I applied the exponent rule and got a square in the bottom? I guess that's the tricky part. After you get a square, you can apply a trig sub. Understand why I chose secant?

OpenStudy (anonymous):

I got to that point. However, I got lost at that part. How do we use sec?

OpenStudy (anonymous):

@sourwig The teacher asks us to not use that method.

zepdrix (zepdrix):

\[\Large\bf\sf \tan^2\theta+1=\sec^2\theta \qquad\to\qquad \color{royalblue}{\sec^2\theta-1=\tan^2 \theta}\]Using secant (and it's square identity) will allow us to get rid of the subtraction under the root. We'll be able to simplify it from there.

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits \frac{1}{\sqrt{\left(\sec \theta\right)^2-1}}dx\quad=\quad \int\limits \frac{1}{\tan \theta}\;dx\]Understand how I did that? I skipped a couple steps in there. We just need to deal with the differential dx before we can integrate.

OpenStudy (anonymous):

I think I got it. But, please, send what you are about to write. It is always useful.

OpenStudy (anonymous):

dx=tanthetadtheta.

OpenStudy (anonymous):

integral of 1 dtheta

OpenStudy (anonymous):

=theta

OpenStudy (anonymous):

=arcsec(e^x)+C

OpenStudy (anonymous):

oh boy oh boy i got this :)

zepdrix (zepdrix):

Oh good c: